# Already Moderate Population Sizes Provably Yield Strong Robustness to Noise

### v3

@@ Abstract
Experience shows that typical evolutionary algorithms can cope well with stochastic disturbances such as noisy function evaluations. In this first mathematical runtime analysis of the $(1+\lambda)$ and $(1,\lambda)$ evolutionary algorithms in the presence of prior bit-wise noise, we show that both algorithms can tolerate constant noise probabilities without increasing the asymptotic runtime on the OneMax benchmark. For this, a population size $\lambda$ suffices that is at least logarithmic in the problem size $n$. The only previous result in this direction regarded the less realistic one-bit noise model, required a population size super-linear in the problem size, and proved a runtime guarantee roughly cubic in the noiseless runtime for the OneMax benchmark. Our significantly stronger results are based on the novel proof argument that the noiseless offspring can be seen as a biased uniform crossover between the parent and the noisy offspring. We are optimistic that the technical lemmas resulting from this insight will find applications also in future mathematical runtime analyses of evolutionary algorithms.
@@ 3.4 Auxiliary Tools
$\binom{n}{k}\geq\mathopen{}\mathclose{{}\left(\frac{n}{k}}\right)^{k}.$
@@
$\mathopen{}\mathclose{{}\left(1-\frac{x}{n}}\right)^{n}\geq e^{-x}-\frac{x^{2}}{2n}.$
@@
For all $x\in\mathopen{}\mathclose{{}\left[0,1}\right]$ and $\lambda>0$ we have
$1-(1-x)^{\lambda}\geq\frac{1}{2}\min\mathopen{}\mathclose{{}\left\{1,\lambda x}\right\}.$
@@ 4 Offspring Distribution
$\frac{\chi}{n}\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)+\frac{q}{n}\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)=\frac{\chi}{n}+\frac{q}{n}-2\cdot\frac{\chi q}{n^{2}}=\frac{1}{n}\mathopen{}\mathclose{{}\left(\chi+q-\frac{2q\chi}{n}}\right).$
@@
$x^{\prime}_{i}=x_{i}$ $\tilde{x}^{\prime}_{i}=x_{i}$ $\frac{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\vphantom{\big{)}}\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)}{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)+\frac{q\chi}{n^{2}}}$ $1-\frac{q\chi}{n^{2}}-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{3}}}\right)$
$x^{\prime}_{i}=x_{i}$ $\tilde{x}^{\prime}_{i}\not=x_{i}$ $\frac{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\frac{q}{n}}{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\frac{q}{n}+\frac{\chi}{n}\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)}$ $\frac{q}{q+\chi}\mathopen{}\mathclose{{}\left(1\pm o(1)}\right)$
$x^{\prime}_{i}\not=x_{i}$ $\tilde{x}^{\prime}_{i}=x_{i}$ $\frac{\frac{\chi q}{n^{2}}}{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)+\frac{q\chi}{n^{2}}}$ $\frac{q\chi}{n^{2}}+O\mathopen{}\mathclose{{}\left(\frac{1}{n^{3}}}\right)$
$x^{\prime}_{i}\not=x_{i}$ $\tilde{x}^{\prime}_{i}\not=x_{i}$ $\frac{\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)\frac{\chi}{n}}{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\frac{q}{n}+\frac{\chi}{n}\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)}$ $\frac{\chi}{q+\chi}\mathopen{}\mathclose{{}\left(1\pm o(1)}\right)$
@@
$\Pr[$ $x_{i}^{\prime}=x_{i}\mid\tilde{x}^{\prime}_{i}=x_{i}]=\frac{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)}{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)+\frac{q\chi}{n^{2}}}$
$=1-\frac{\frac{q\chi}{n^{2}}}{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)+\frac{q\chi}{n^{2}}}=1-\frac{q\chi}{n^{2}}\cdot\frac{1}{\mathopen{}\mathclose{{}\left(1-O\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)}\right)}$
$=1-\frac{q\chi}{n^{2}}-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{3}}}\right).$
@@
$\Pr[$ $x_{i}^{\prime}=x_{i}\mid\tilde{x}^{\prime}_{i}\not=x_{i}]=\frac{\Pr[x_{i}^{\prime}=x_{i}\not=\tilde{x}^{\prime}_{i}]}{\Pr[x_{i}\not=\tilde{x}^{\prime}_{i}]}=\frac{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\frac{q}{n}}{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\frac{q}{n}+\frac{\chi}{n}\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)}$
$=\frac{q}{q+\chi\mathopen{}\mathclose{{}\left(\frac{1-\frac{q}{n}}{1-\frac{\chi}{n}}}\right)}=\frac{q}{q+\chi}\cdot\frac{q+\chi}{q+\chi\pm o\mathopen{}\mathclose{{}\left(1}\right)}=\frac{q}{q+\chi}\mathopen{}\mathclose{{}\left(1\pm o\mathopen{}\mathclose{{}\left(1}\right)}\right),$
@@
$=\frac{\frac{\chi q}{n^{2}}}{\mathopen{}\mathclose{{}\left(1-\frac{\chi}{n}}\right)\mathopen{}\mathclose{{}\left(1-\frac{q}{n}}\right)+\frac{q\chi}{n^{2}}}=\frac{q\chi}{n^{2}}\mathopen{}\mathclose{{}\left(1+O\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)}\right)=\frac{q\chi}{n^{2}}+O\mathopen{}\mathclose{{}\left(\frac{1}{n^{3}}}\right),$
@@
$=\frac{\chi}{q+\chi}\mathopen{}\mathclose{{}\left(1\pm o\mathopen{}\mathclose{{}\left(1}\right)}\right).\qed$
@@
$\geq\mathopen{}\mathclose{{}\left(d(x)-d(z)}\right)\cdot\frac{\chi}{q+\chi}\cdot(1\pm o(1))-\frac{q\chi}{n}-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{2}}}\right).$
@@
$=(D+C)+(A-D)\mathopen{}\mathclose{{}\left(\frac{q\chi}{n^{2}}+O\mathopen{}\mathclose{{}\left(\frac{1}{n^{3}}}\right)}\right)+(B-C)\frac{\chi}{q+\chi}(1\pm o(1))$
$\leq d(x)+n\mathopen{}\mathclose{{}\left(\frac{q\chi}{n^{2}}+O\mathopen{}\mathclose{{}\left(\frac{1}{n^{3}}}\right)}\right)+\mathopen{}\mathclose{{}\left(d(z)-d(x)}\right)\frac{\chi}{q+\chi}(1\pm o(1))$
$=d(x)+\frac{q\chi}{n}+O\mathopen{}\mathclose{{}\left(\frac{1}{n^{2}}}\right)+\mathopen{}\mathclose{{}\left(d(z)-d(x)}\right)\frac{\chi}{q+\chi}(1\pm o(1)),$
@@ 5 Runtime Analysis
We use the results from Section 4 to estimate upper bounds on the runtime of the $(1+\lambda)$ EA and the $(1,\lambda)$ EA. The main result of this section is the following theorem.
@@
Consider a run of the $(1,\lambda)$ EA or the $(1+\lambda)$ EA with mutation rate $\frac{\chi}{n}$ where $\chi=\Theta(1)$ on the OneMax problem with bit-wise noise with rate $\frac{q}{n}$ where $q=O(1)$. If the population size $\lambda$ is at least $C\ln(n)$ for some constant $C$ depending on $\chi$ and $q$, then the expected number $E[T_{F}]$ of fitness evaluations until the optimum is sampled is
$O\mathopen{}\mathclose{{}\left(n\log(n)+n\lambda\frac{\log\log(\lambda)}{\log(\lambda)}}\right).$
To prove Theorem 7, we use the variable drift theorem (Theorem 1). The process we apply it to is $\{d_{t}\}_{t\in{\mathbb{N}}_{0}}$, which is the distance of the current individual (of either the $(1,\lambda)$ EA or the $(1+\lambda)$ EA) to the optimum after iteration $t$. Since we consider the OneMax function, $d_{t}$ is equal to the number of zero-bits in that individual.
@@
$\Delta^{\text{com}}(d)$ $\geq(1-o(1))\frac{\chi}{q+\chi}\mathopen{}\mathclose{{}\left(\Delta^{+}(d)-\Delta^{-}(d)}\right)-\frac{q\chi}{n}-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{2}}}\right)$ (1)
@@
$\begin{split}\Delta^{\text{plus}}(d)&\geq(1-o(1))\frac{\chi}{q+\chi}\mathopen{}\mathclose{{}\left((1-o(1))e^{-q}\Delta^{+}(d)-\Delta^{-}(d)}\right)\\ &\quad\,-\frac{q\chi}{n}-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{2}}}\right)\end{split}$ (2)
@@
$\Delta^{\text{com}}(d)$
$\geq\sum_{z\in\{0,1\}^{n}}\Pr[\tilde{y}=z]\bigg{(}\mathopen{}\mathclose{{}\left(d-d_{z}}\right)\frac{\chi}{q+\chi}(1\pm o(1))-\frac{q\chi}{n}-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{2}}}\right)\bigg{)}$
$=\mathopen{}\mathclose{{}\left((1-o(1))\frac{\chi}{q+\chi}\sum_{z\in\{0,1\}^{n}}\Pr[\tilde{y}=z]\mathopen{}\mathclose{{}\left(d-d_{z}}\right)}\right)-\frac{q\chi}{n}-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{2}}}\right)$
$=(1-o(1))\frac{\chi}{q+\chi}E[d-\tilde{d}_{y}]-\frac{q\chi}{n}-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{2}}}\right).$
@@
$d-d_{\text{plus}}=(d-d_{y})\mathds{I}\mathopen{}\mathclose{{}\left[\tilde{d}\geq\tilde{d}_{y}}\right],$ (3)
@@
$\Delta$ (${}^{\text{plus}}(d)=E[d-d_{\text{plus}}]=E\mathopen{}\mathclose{{}\left[(d-d_{y})\mathds{I}\mathopen{}\mathclose{{}\left[\tilde{d}\geq\tilde{d}_{y}}\right]}\right]$
$=\sum_{z\in\{0,1\}^{n}}\Pr[\tilde{y}=z]\cdot\bigg{(}\Pr\mathopen{}\mathclose{{}\left[\tilde{d}\geq\tilde{d}_{y}\mid\tilde{y}=z}\right]$
$\quad\,\cdot E\mathopen{}\mathclose{{}\left[d-d_{y}\mid\tilde{y}=z,\tilde{d}\geq\tilde{d}_{y}}\right]+\Pr\mathopen{}\mathclose{{}\left[\tilde{d}<\tilde{d}_{y}\mid\tilde{y}=z}\right]\cdot 0\bigg{)}.$
@@
$E\Big{[}d-d_{y}$ $\mid\tilde{y}=z,\tilde{d}\geq\tilde{d}_{y}\Big{]}=E\mathopen{}\mathclose{{}\left[d-d_{y}\mid\tilde{y}=z}\right]$
$\geq(1\pm o(1))\frac{\chi}{q+\chi}\mathopen{}\mathclose{{}\left(d-d_{z}}\right)-\frac{q\chi}{n}-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{2}}}\right).$
@@
$\begin{split}\Delta^{\text{plus}}(d)&\geq(1-o(1))\frac{\chi}{q+\chi}\\ &\quad\,\cdot\sum_{z\in\{0,1\}^{n}}\Pr[\tilde{y}=z]\Pr[\tilde{d}\geq\tilde{d}_{y}\mid\tilde{y}=z](d-d_{z})\\ &\quad\,-\frac{q\chi}{n}-O\mathopen{}\mathclose{{}\left(\frac{1}{n^{2}}}\right).\end{split}$ (4)
@@
Above we obtained a weaker drift estimate for the $(1+\lambda)$ EA than the $(1,\lambda)$ EA. This is counter-intuitive – one would feel that the $(1+\lambda)$ EA should profit to some extent from the property that the current-best solution is participating in the selection of the next parent. The reason for our weaker bound in the estimate following eq. 4, where we pessimistically estimated that inferior mutation winners $y$ are always accepted, but superior ones only when the parent is not subject to noise. We feel that this cannot be avoided in the general case (note that Lemma 8 is valid in general and is not specific for the optimization of OneMax). We are sure that when exploiting properties of OneMax, stronger bounds could be shown. We refrain from this both because we like our general, problem-independent approach and because all we can gain are constant factors, which are generally ignored in an asymptotic analysis as ours (note that we lose constant factors and lower order terms in the later part of the analysis anyway).
In the following lemmas we show estimates for $\Delta^{+}(d)$ and $\Delta^{-}(d)$, which naturally are specific to the OneMax problem. Lemmas 9-11 estimate the positive component of the drift for different distances. We note that implicitly these results exist in the proofs in [DK15], but distilling them from there is rather complicated.
@@
If $\lambda=\omega(1)$ then for any distance $d\geq\frac{n}{\ln\lambda}$ we have $\Pr[d-\tilde{d}_{y}\geq\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor]\geq\frac{1}{4}$ and the positive component of the drift is at least
$\Delta^{+}(d)\geq\frac{1}{4}\mathopen{}\mathclose{{}\left\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}}\right\rfloor.$
@@
Let $x^{\prime}$ be some arbitrary offspring and let $\tilde{x}^{\prime}$ be the same offspring after noise. Denote the distances from them to the optimum by $d_{x^{\prime}}$ and $\tilde{d}_{x^{\prime}}$ correspondingly. Let also $p_{i}=\Pr[d-\tilde{d}_{x^{\prime}}\geq i]\geq\Pr[d-\tilde{d}_{x^{\prime}}=i]\geq\binom{d}{i}\mathopen{}\mathclose{{}\left(\frac{r}{n}}\right)^{i}\mathopen{}\mathclose{{}\left(1-\frac{r}{n}}\right)^{n-i}$, since each noisy offspring is distributed as if it was created via standard bit mutation with rate $\frac{r}{n}$ where $r=\chi+q-\frac{2q\chi}{n}$. Then we compute
$p_{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}$ $\geq\binom{d}{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\mathopen{}\mathclose{{}\left(\frac{r}{n}}\right)^{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\mathopen{}\mathclose{{}\left(1-\frac{r}{n}}\right)^{n-\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}$
$\geq\mathopen{}\mathclose{{}\left(\frac{d}{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}}\right)^{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\mathopen{}\mathclose{{}\left(\frac{r}{n}}\right)^{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\mathopen{}\mathclose{{}\left(1-\frac{r}{n}}\right)^{n}$
$\geq\mathopen{}\mathclose{{}\left(\frac{dr}{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor n}}\right)^{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\mathopen{}\mathclose{{}\left(e^{-r}-\frac{r^{2}}{2n}}\right),$
@@
$p_{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}$ $\geq\exp\mathopen{}\mathclose{{}\left(\frac{\ln\lambda}{2\ln\ln\lambda}\mathopen{}\mathclose{{}\left(\ln{\frac{dr}{n}}-\ln{\frac{\ln\lambda}{2\ln\ln\lambda}}}\right)-2r}\right)$
$\geq\exp\bigg{(}\frac{\ln\lambda}{2\ln\ln\lambda}\bigg{(}\ln\frac{r}{\ln\lambda}-\ln\ln\lambda+\ln(2\ln\ln\lambda)$
$-2r\cdot\frac{2\ln\ln\lambda}{\ln\lambda}\bigg{)}\bigg{)}$
@@
$p_{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\geq\exp\mathopen{}\mathclose{{}\left(\frac{\ln\lambda}{2\ln\ln\lambda}\cdot(-2\ln\ln\lambda)}\right)=e^{-\ln\lambda}=\frac{1}{\lambda}.$

### v4

@@ Abstract
Experience shows that typical evolutionary algorithms can cope well with stochastic disturbances such as noisy function evaluations. In this first mathematical runtime analysis of the $(1+\lambda)$ and $(1,\lambda)$ evolutionary algorithms in the presence of prior bit-wise noise, we show that both algorithms can tolerate constant noise probabilities without increasing the asymptotic runtime on the OneMax benchmark. For this, a population size $\lambda$ suffices that is at least logarithmic in the problem size $n$. The only previous result in this direction regarded the less realistic one-bit noise model, required a population size super-linear in the problem size, and proved a runtime guarantee roughly cubic in the noiseless runtime for the OneMax benchmark. Our significantly stronger results are based on the novel proof argument that the noiseless offspring can be seen as a biased uniform crossover between the parent and the noisy offspring. We are optimistic that this technique will find applications also in future mathematical runtime analyses of evolutionary algorithms.
@@ 3.4 Auxiliary Tools
$\binom{n}{k}\geq\left(\frac{n}{k}\right)^{k}.$
@@
$\left(1-\frac{x}{n}\right)^{n}\geq e^{-x}-\frac{x^{2}}{2n}.$
@@
For all $x\in\left[0,1\right]$ and $\lambda>0$ we have
$1-(1-x)^{\lambda}\geq\frac{1}{2}\min\left\{1,\lambda x\right\}.$
@@ 4 Offspring Distribution
$\frac{\chi}{n}\left(1-\frac{q}{n}\right)+\frac{q}{n}\left(1-\frac{\chi}{n}\right)=\frac{\chi}{n}+\frac{q}{n}-2\cdot\frac{\chi q}{n^{2}}=\frac{1}{n}\left(\chi+q-\frac{2q\chi}{n}\right).$
@@
$x^{\prime}_{i}=x_{i}$ $\tilde{x}^{\prime}_{i}=x_{i}$ $\frac{\left(1-\frac{\chi}{n}\right)\vphantom{\big{)}}\left(1-\frac{q}{n}\right)}{\left(1-\frac{\chi}{n}\right)\left(1-\frac{q}{n}\right)+\frac{q\chi}{n^{2}}}$ $1-\frac{q\chi}{n^{2}}-O\left(\frac{1}{n^{3}}\right)$
$x^{\prime}_{i}=x_{i}$ $\tilde{x}^{\prime}_{i}\not=x_{i}$ $\frac{\left(1-\frac{\chi}{n}\right)\frac{q}{n}}{\left(1-\frac{\chi}{n}\right)\frac{q}{n}+\frac{\chi}{n}\left(1-\frac{q}{n}\right)}$ $\frac{q}{q+\chi}\left(1\pm o(1)\right)$
$x^{\prime}_{i}\not=x_{i}$ $\tilde{x}^{\prime}_{i}=x_{i}$ $\frac{\frac{\chi q}{n^{2}}}{\left(1-\frac{\chi}{n}\right)\left(1-\frac{q}{n}\right)+\frac{q\chi}{n^{2}}}$ $\frac{q\chi}{n^{2}}+O\left(\frac{1}{n^{3}}\right)$
$x^{\prime}_{i}\not=x_{i}$ $\tilde{x}^{\prime}_{i}\not=x_{i}$ $\frac{\left(1-\frac{q}{n}\right)\frac{\chi}{n}}{\left(1-\frac{\chi}{n}\right)\frac{q}{n}+\frac{\chi}{n}\left(1-\frac{q}{n}\right)}$ $\frac{\chi}{q+\chi}\left(1\pm o(1)\right)$
@@
$\Pr[$ $x_{i}^{\prime}=x_{i}\mid\tilde{x}^{\prime}_{i}=x_{i}]=\frac{\left(1-\frac{\chi}{n}\right)\left(1-\frac{q}{n}\right)}{\left(1-\frac{\chi}{n}\right)\left(1-\frac{q}{n}\right)+\frac{q\chi}{n^{2}}}$
$=1-\frac{\frac{q\chi}{n^{2}}}{\left(1-\frac{\chi}{n}\right)\left(1-\frac{q}{n}\right)+\frac{q\chi}{n^{2}}}=1-\frac{q\chi}{n^{2}}\cdot\frac{1}{\left(1-O\left(\frac{1}{n}\right)\right)}$
$=1-\frac{q\chi}{n^{2}}-O\left(\frac{1}{n^{3}}\right).$
@@
$\Pr[$ $x_{i}^{\prime}=x_{i}\mid\tilde{x}^{\prime}_{i}\not=x_{i}]=\frac{\Pr[x_{i}^{\prime}=x_{i}\not=\tilde{x}^{\prime}_{i}]}{\Pr[x_{i}\not=\tilde{x}^{\prime}_{i}]}=\frac{\left(1-\frac{\chi}{n}\right)\frac{q}{n}}{\left(1-\frac{\chi}{n}\right)\frac{q}{n}+\frac{\chi}{n}\left(1-\frac{q}{n}\right)}$
$=\frac{q}{q+\chi\left(\frac{1-\frac{q}{n}}{1-\frac{\chi}{n}}\right)}=\frac{q}{q+\chi}\cdot\frac{q+\chi}{q+\chi\pm o\left(1\right)}=\frac{q}{q+\chi}\left(1\pm o\left(1\right)\right),$
@@
$=\frac{\frac{\chi q}{n^{2}}}{\left(1-\frac{\chi}{n}\right)\left(1-\frac{q}{n}\right)+\frac{q\chi}{n^{2}}}=\frac{q\chi}{n^{2}}\left(1+O\left(\frac{1}{n}\right)\right)=\frac{q\chi}{n^{2}}+O\left(\frac{1}{n^{3}}\right),$
@@
$=\frac{\chi}{q+\chi}\left(1\pm o\left(1\right)\right).\qed$
@@
$\geq\left(d(x)-d(z)\right)\cdot\frac{\chi}{q+\chi}\cdot(1\pm o(1))-\frac{q\chi}{n}-O\left(\frac{1}{n^{2}}\right).$
@@
$=(D+C)+(A-D)\left(\frac{q\chi}{n^{2}}+O\left(\frac{1}{n^{3}}\right)\right)+(B-C)\frac{\chi}{q+\chi}(1\pm o(1))$
$\leq d(x)+n\left(\frac{q\chi}{n^{2}}+O\left(\frac{1}{n^{3}}\right)\right)+\left(d(z)-d(x)\right)\frac{\chi}{q+\chi}(1\pm o(1))$
$=d(x)+\frac{q\chi}{n}+O\left(\frac{1}{n^{2}}\right)+\left(d(z)-d(x)\right)\frac{\chi}{q+\chi}(1\pm o(1)),$
@@ 5 Runtime Analysis
We use the results from Section 4 to prove upper bounds on the runtime of the $(1+\lambda)$ EA and the $(1,\lambda)$ EA. The main result of this section is the following theorem.
@@
Consider a run of the $(1,\lambda)$ EA or the $(1+\lambda)$ EA with mutation rate $\frac{\chi}{n}$ on the OneMax problem with bit-wise noise with rate $\frac{q}{n}$. Here $\chi$ can be any positive constant (independent of the problem size $n$). The parameter $q$ may depend on $n$ as long as $q=O(1)$. If the population size $\lambda$ is at least $C\ln(n)$ for some constant $C$ depending on $\chi$ and $q$, then the expected number $E[T_{F}]$ of fitness evaluations until the optimum is sampled is
$O\left(n\log(n)+n\lambda\frac{\log\log(\lambda)}{\log(\lambda)}\right).$
To prove Theorem 7, we use the variable drift theorem (Theorem 1). The process we apply it to is $\{d_{t}\}_{t\in{\mathbb{N}}_{0}}$, which is the distance of the current individual (of either the $(1,\lambda)$ EA or the $(1+\lambda)$ EA) to the optimum after iteration $t$. Since we consider the OneMax problem, $d_{t}$ is equal to the number of zero-bits in that individual.
@@
$\Delta^{\text{com}}(d)$ $\geq(1-o(1))\frac{\chi}{q+\chi}\left(\Delta^{+}(d)-\Delta^{-}(d)\right)-\frac{q\chi}{n}-O\left(\frac{1}{n^{2}}\right)$ (1)
@@
$\begin{split}\Delta^{\text{plus}}(d)&\geq(1-o(1))\frac{\chi}{q+\chi}\left((1-o(1))e^{-q}\Delta^{+}(d)-\Delta^{-}(d)\right)-\frac{q\chi}{n}-O\left(\frac{1}{n^{2}}\right)\end{split}$ (2)
@@
$\Delta^{\text{com}}(d)$ $\geq\sum_{z\in\{0,1\}^{n}}\Pr[\tilde{y}=z]\bigg{(}\left(d-d_{z}\right)\frac{\chi}{q+\chi}(1\pm o(1))-\frac{q\chi}{n}-O\left(\frac{1}{n^{2}}\right)\bigg{)}$
$=\left((1-o(1))\frac{\chi}{q+\chi}\sum_{z\in\{0,1\}^{n}}\Pr[\tilde{y}=z]\left(d-d_{z}\right)\right)-\frac{q\chi}{n}-O\left(\frac{1}{n^{2}}\right)$
$=(1-o(1))\frac{\chi}{q+\chi}E[d-\tilde{d}_{y}]-\frac{q\chi}{n}-O\left(\frac{1}{n^{2}}\right).$
@@
$d-d_{\text{plus}}=(d-d_{y})\mathds{I}\left[\tilde{d}\geq\tilde{d}_{y}\right],$ (3)
@@
$\Delta$ (${}^{\text{plus}}(d)=E[d-d_{\text{plus}}]=E\left[(d-d_{y})\mathds{I}\left[\tilde{d}\geq\tilde{d}_{y}\right]\right]$
$=\sum_{z\in\{0,1\}^{n}}\Pr[\tilde{y}=z]\cdot\bigg{(}\Pr\left[\tilde{d}\geq\tilde{d}_{y}\mid\tilde{y}=z\right]$
$\quad\,\cdot E\left[d-d_{y}\mid\tilde{y}=z,\tilde{d}\geq\tilde{d}_{y}\right]+\Pr\left[\tilde{d}<\tilde{d}_{y}\mid\tilde{y}=z\right]\cdot 0\bigg{)}.$
@@
$E\Big{[}d-d_{y}$ $\mid\tilde{y}=z,\tilde{d}\geq\tilde{d}_{y}\Big{]}=E\left[d-d_{y}\mid\tilde{y}=z\right]$
$\geq(1\pm o(1))\frac{\chi}{q+\chi}\left(d-d_{z}\right)-\frac{q\chi}{n}-O\left(\frac{1}{n^{2}}\right).$
@@
$\begin{split}\Delta^{\text{plus}}(d)&\geq(1-o(1))\frac{\chi}{q+\chi}\cdot\sum_{z\in\{0,1\}^{n}}\Pr[\tilde{y}=z]\Pr[\tilde{d}\geq\tilde{d}_{y}\mid\tilde{y}=z](d-d_{z})\\ &\quad\,-\frac{q\chi}{n}-O\left(\frac{1}{n^{2}}\right).\end{split}$ (4)
@@
Above we obtained a weaker drift estimate for the $(1+\lambda)$ EA than the $(1,\lambda)$ EA. This is counter-intuitive – one would feel that the $(1+\lambda)$ EA should profit to some extent from the property that the current-best solution is participating in the selection of the next parent. The reason for our weaker bound is the estimate following eq. (4), where we pessimistically estimated that inferior mutation winners $y$ are always accepted, but superior ones only when the parent is not subject to noise. We feel that this cannot be avoided in the general case (note that Lemma 8 is valid in general and is not specific for the optimization of OneMax). We are sure that when exploiting properties of OneMax, stronger bounds could be shown. We refrain from this both because we like our general, problem-independent approach and because all we can gain are constant factors, which are generally ignored in an asymptotic analysis as ours (note that we lose constant factors and lower order terms in the later part of the analysis anyway).
In the following lemmas we show estimates for $\Delta^{+}(d)$ and $\Delta^{-}(d)$, which naturally are specific to the OneMax problem. Lemmas 9 to 11 estimate the positive component of the drift for different distances. We note that implicitly these results exist in the proofs in [DK15], but distilling them from there is rather complicated.
@@
If $\lambda=\omega(1)$, then for any distance $d\geq\frac{n}{\ln\lambda}$ we have $\Pr[d-\tilde{d}_{y}\geq\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor]\geq\frac{1}{4}$ and the positive component of the drift is at least
$\Delta^{+}(d)\geq\frac{1}{4}\left\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\right\rfloor.$
@@
Let $x^{\prime}$ be some arbitrary offspring and let $\tilde{x}^{\prime}$ be the same offspring after noise. Denote the distances from them to the optimum by $d_{x^{\prime}}$ and $\tilde{d}_{x^{\prime}}$ correspondingly. Let also $p_{i}=\Pr[d-\tilde{d}_{x^{\prime}}\geq i]\geq\Pr[d-\tilde{d}_{x^{\prime}}=i]\geq\binom{d}{i}\left(\frac{r}{n}\right)^{i}\left(1-\frac{r}{n}\right)^{n-i}$, since each noisy offspring is distributed as if it was created via standard bit mutation with rate $\frac{r}{n}$ where $r=\chi+q-\frac{2q\chi}{n}$. Then we compute
$p_{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}$ $\geq\binom{d}{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\left(\frac{r}{n}\right)^{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\left(1-\frac{r}{n}\right)^{n-\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}$
$\geq\left(\frac{d}{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\right)^{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\left(\frac{r}{n}\right)^{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\left(1-\frac{r}{n}\right)^{n}$
$\geq\left(\frac{dr}{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor n}\right)^{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\left(e^{-r}-\frac{r^{2}}{2n}\right),$
@@
$p_{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}$ $\geq\exp\left(\frac{\ln\lambda}{2\ln\ln\lambda}\left(\ln{\frac{dr}{n}}-\ln{\frac{\ln\lambda}{2\ln\ln\lambda}}\right)-2r\right)$
$\geq\exp\bigg{(}\frac{\ln\lambda}{2\ln\ln\lambda}\bigg{(}\ln\frac{r}{\ln\lambda}-\ln\ln\lambda+\ln(2\ln\ln\lambda)-2r\cdot\frac{2\ln\ln\lambda}{\ln\lambda}\bigg{)}\bigg{)}.$
@@
$p_{\lfloor\frac{\ln\lambda}{2\ln\ln\lambda}\rfloor}\geq\exp\left(\frac{\ln\lambda}{2\ln\ln\lambda}\cdot(-2\ln\ln\lambda)\right)=e^{-\ln\lambda}=\frac{1}{\lambda}.$